In an isosceles triangle, one angle is 12 more than twice the measure of one of the equal angles. What is the measure of each angle?
Answer by mananth(16949)

let equal angle be x
the third angle = 2x+12
....
x+x+2x+12=180
4x+12=180
4x=180-12
4x=168
/4
x=168/4
x= 42 deg
42,42, 96